OpenStudy (anonymous):

solve the following equation for y -27x+3y=-6

OpenStudy (anonymous):

Is there an another equation?

OpenStudy (anonymous):

no its just that

OpenStudy (anonymous):

OK then, 3y = -6 + 27x y = (-6 + 27x)/3 y = -2 + 9x

OpenStudy (anonymous):

y=9x−2

OpenStudy (anonymous):

the answers y=9x-2 ?

OpenStudy (anonymous):

everyone give each other medals i can't give all of you medals

OpenStudy (anonymous):

@jaredhm29

OpenStudy (anonymous):

yes if u need a good calc online search math papa algebra calc put in the equation and it gives u step by step

OpenStudy (anonymous):

@NadiaElhaj_1

OpenStudy (anonymous):

\[-27x+3y=-6\\ \implies\ -9x+y=-2\\ \implies y=9x+2\\ now,\ put\ x= any \ number\ get \ corresponding\ value\ of\ y.\\this\ equation\ has\ infinite\ solutions.\]

OpenStudy (anonymous):

thank you guys so much medal each other for me please

OpenStudy (anonymous):

np

OpenStudy (anonymous):

medals doesn't matter, you get solutions. that's enough!